WLCG-OSG-EGEE Operations meeting

Europe/Zurich
28-R-15 (CERN conferencing service (joining details below))

28-R-15

CERN conferencing service (joining details below)

Description
grid-operations-meeting@cern.ch
Weekly OSG, EGEE, WLCG infrastructure coordination meeting.
We discuss the weekly running of the production grid infrastructure based on weekly reports from the attendees. The reported issues are discussed, assigned to the relevant teams, followed up and escalated when needed. The meeting is also the forum for the sites to get a summary of the weekly WLCG activities and plans
Attendees:
  • OSG operations team
  • EGEE operations team
  • EGEE ROC managers
  • WLCG coordination representatives
  • WLCG Tier-1 representatives
  • other site representatives (optional)
  • GGUS representatives
  • VO representatives
  • To dial in to the conference:
    a. Dial +41227676000
    b. Enter access code 0140768

    OR click HERE

    NB: Reports were not received in advance of the meeting from:

  • ROCs: CERN
  • VOs:
  • <big> Items from Alice </big>
  • <big> Items from ATLAS </big>
  • <big> Items from CMS </big>
    • On of the main focus of last week was T1 workflows. The first was stable, some Castor issues (mostly GC-related) addressed and fixed. On the second, we had reprocessing and skimming jobs running at T1 sites.
      ASGC: Pretty impressive performances; no problems found.
      CNAF also OK, running up to ~600 skim jobs in parallel.
      FNAL is running all kind of processing jobs since days, up to 3.4K running in parallel.
      FZK running processing with some issues reading the input RAW data, skim jobs also somehow slower than other T1's.
      IN2P3 is clearing out its backlog, more processing jobs will come.
      PIC has much processing to go, total number of slots used driven by the number of running skim jobs, some issues there are being investigated with the input/help from other T1's.
      RAL using all queues, some jobs killed by some CEs because exceeding the maximum CPU time on them, now restricted to use the long queue, and things got better.
    • Interesting week with transfers, especially in T0-T1 together with other VO's, and in T1-T2. Data transfer tests stable in the T0-T1 routes, and will continue until the end of the challenge. T1-T2 going on with link rotation, details at https://twiki.cern.ch/twiki/bin/view/CMS/CCRC08-DataTransfers . Tests ramping down in T1-T1 routes.
    • at the T0, repacker tests in progress. More info will be collected during this week.
    Speaker: Daniele Bonacorsi
  • <big> Items from LHCb </big>
  • 5:00 PM 5:30 PM
    OSG Items 30m
    None received.
    Speaker: Rob Quick (OSG - Indiana University)
  • 5:30 PM 5:35 PM
    Review of action items 5m
    list of actions
  • 5:35 PM 5:35 PM
    AOB